Solve the valid solutions of the following equations providing your values for some of them. Write your answers in solution set form. Identify the extraneous solution, if there is.


a^x-1 = b^3-x


(provide values for a and b which are different integers greater than 1)


1
Expert's answer
2021-05-19T17:47:04-0400

"a^{x-1}=b^{3-x}"

"\ln a*(x-1)=\ln b* (3-x)"

"x=\frac{3\ln a+\ln b}{\ln a+\\ln b}"

For a=2, b=4:

"x=\frac{3\ln 2+\ln 4}{\ln 2+\\ln 4}=\frac{5\ln 2}{3\ln 2}=\frac{5}{3}"